OpenStudy (anonymous):

A model rocket is launched from point A with an initial velocity v of 86 m/s. If the rocket's descent parachute does not deploy and the rocket lands 104m from A, determine (a)the angle that v forms with the vertical, (b)the maximum height h reached by the rocket, and (c)the duration of the flight.

OpenStudy (whpalmer4):

Height of the rocket will be \[h(t) = -\frac{1}{2}gt^2 + v_0 t\sin \theta + h_0\] where \(\qquad g=9.8 \text{ m/s}^2\) \(\qquad v_0 = 86 \text{ m/s}\) \(\qquad h_0 = 0 \text{ m}\) \(\qquad \theta = \text{ angle formed with the vertical}\) That's a parabola. You'll solve that for \(h(t_f) = 0\) to find the time of flight. The horizontal component of the rocket's velocity will be \(v_x = v_0 \cos \theta\). You know that \(x = v_x t_f = 104 \text{ m}\) where \(t_f\) is the time of flight. You can use that relationship to write an expression for \(t_f\) in terms of \(v_0\) and \(\theta\). Substitute that into the first equation and solve for \(\theta\). Once you've got the parabola figured out, you can easily find the maximum height by finding the vertex, and you've already found the duration of the flight.